Electricar Inc. is considering expanding manufacturing capacities in Asia. This would involve an initial investment of 4 billion RMB . The plant will start production after one year. It is expected to last for five years and a have a salvage value and the end of this period of 500 million RMB in real terms. The plant will produce 100 000 cars a year. The firm anticipates that in the first year it will be able to sell each car for 65 000 RMB, and thereafter the price is expected to increase by 4% a year.

Raw materials for each car are forecasted to cost 20 000 RMB in the first year and these costs are predicted to increase by 3% yearly. Total labor costs for the plant are expected to be 1.1 billion RMB in the first year and thereafter will increase by 7% a year. The land on which the plant is built can be rented for five years at a fixed cost of 300 million RMB a year payable at the start of each year. Electricar Inc.'s discount rate for this type of project is 10% (nominal). The expected rate of inflation is 5%. The plant can be depreciated straight-line over the five-year period and profits will be taxed at 20%. Assume all cash flows occur at the end of each except where otherwise stated.

What is the NPV of the project plant for Electricar Inc?
Answer #1

We calculate the nominal cash flows for each year, and discount them at the nominal rate to the find the NPV in real terms.

Nominal salvage value of plant after 5 years = real salvage value * (1 + inflation rate)5

Nominal salvage value of plant after 5 years = 500,000,000 * (1 + 5%)5 = RMB 638,140,781

Annual depreciation = (Cost of plant - salvage value) / depreicable life

Annual depreciation = (4,000,000,000 - 638,140,781) / 5 = RMB 672,371,844

Operating cash flow (OCF) each year = income after tax + depreciation

The salvage value equals the book value of the plant, and hence there is no tax adjustment required.

NPV is calculated using NPV function in Excel

NPV is RMB 6,908,086,502

ei ol 2 Initial Investment 3 Cost of plant 4,000,000,000 100,000 76,041 22,510 7,604,080,640 2,251,017,620 5 OCF 6 Number of

A в G 2 2 Initial Investment 3 Cost of plant 4000000000 100000 =F7*(1+4%) =F8*(1+3%) =G6 G7 =G6-G8 300000000 5 OCF 6 Number o